# Privacy in Speech Technology

### v3

@@ VI Evaluating the performance of privacy-preserving methods
To evaluate the performance of any privacy-preserving methods, we need performance measures for both utility and privacy, corresponding respectively to the trusted and threat tasks (see fig. 2). With utility, we refer here to the usefulness of the service in its intended task. While the objective performance in both utility and privacy are prerequisites for any good speech application, to the user, it is central that their experience with the service or application is good. Correspondingly, performance can always be measured on two dimensions, privacy and utility, as well as objective and subjective (see table IV). Performance with respect to objective and subjective privacy and utility, are presented in the following subsections (sections VI-A, VI-B, VI-C and VI-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VII). We further need to define how the objective measures are applied in systems design to ensure privacy (see section V-F).
The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [98]. It is based on direct measurement of identifiability using signal processing and machine learning methods. The evaluation objective in such empirical privacy is to determine the extent to which an attacker can re-identify an anonymized speaker or anonymized attribute(s) of a speaker (see fig. 3).
@@ VI-A Objective measures of privacy
A central issue with these metrics is that they are based on mathematical models of probability distributions, whereas speech is a natural signal that does not submit to exact formulas [87]. To use these formulas, we must, therefore, first choose models to describe the distributions of observations and then estimate their parameters. Both steps lead to inaccuracies that dilute the evidentiary power of the proofs of privacy.
Measures commonly used to analyze privacy in speaker recognition, -verification, and biometrics include [172]:
@@
    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [56, 88]. An increase in EER means that the attacker has made more errors, and privacy has improved.
@@
    The log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [173]
@@
    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [56, 174].
@@
    The expected privacy disclosure from the Privacy ZEBRA framework [87] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
@@
    The worst-case disclosure also from the Privacy ZEBRA framework [87] is the largest disclosure observed in a dataset.
@@
As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [98]. Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [35]. The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [123]. More comprehensive treatments of objective privacy metrics are provided in [13, 175].
@@ VI-B Subjective evaluation of privacy
In speaker recognition, computers are now on par or better than humans [36] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [176, 177]. We can thus expect that computers are better at evaluating privacy in well-defined tasks. However, there are at least two caveats. First, in identifying private information, human listeners can potentially use properties or features of speech signals other than those computers use. Subjective listeners can thus potentially discover new weaknesses that computers miss (though the opposite is also true). Second, in scenarios where humans are the potential attackers extracting information (as in threats 3, 5, and 7 in fig. 1), the true measure of privacy is quantified by subjective listening. Thus, it is often prudent to complement objective privacy metrics with subjective tests.
@@
As an example, in the VoicePrivacy 2020 challenge, subjective speaker verifiability was measured by asking subjects to rate, on a scale of 1 to 10110, the similarity between an anonymized target and a reference speaker, where the reference was either the same or a different speaker [98]. If the distributions of similarity scores have significant differences, then the anonymization has failed. Alternatives to the 1 to 10110 Likert-scale include, for example, AB (are these samples the same person?) and ABX (which of AB is equal to X?) tests. For a comprehensive exposition of subjective evaluation methods, see [178].
@@ VI-C Objective performance in downstream tasks
Metrics used for quantifying utility depend on the application. In applications where a computer analyzes the speech signal, we quantify utility by measuring the accuracy of analysis. For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [179]. By measuring the word error rate with and without privacy-preserving processing, we can determine how much privacy-protections impede utility.
@@ VI-D Subjective performance in downstream tasks
If a human user is to listen to the output signal, then that user is the ultimate measure of performance [178]. For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [180]. Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [181, 182]. Similarly, if privacy protections degrade a speech signal, we can readily quantify such degradations by listening tests.
Subjective listening tests are, however, expensive, time-consuming, and noisy. Individual subjects, especially those without training (known as naïve subjects), are typically unable to repeat a sequence of tests such that results would match. Subjects are also slow since they need to listen to samples one at a time, at real-time speeds, require time to answer between each sample, a multitude of subjects need to do the same test to gain statistical confidence in results, and tests need oversight by trained personnel. Automated tests, known as objective measures of perceptual quality, are an alternative to such subjective tests. They estimate the outcome of subjective listening tests by computational means. For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [183]. Later improvements include POLQA and ViSQOL [184, 185]. Such objective measures can be used to automate quality testing at a lower price and faster speed so that repeated estimates with the same input give the same result. Observe that researchers should carefully review the specifications of such algorithms to determine whether they apply to the problem at hand to ensure reliable results.
@@ VII Perception and psychology of privacy
Much of the work of the speech community is focused on technical aspects of privacy-preserving algorithms. Yet speech devices are used by and designed for people, and they use people’s voice data. It is therefore essential to know what people think of speech devices, how they use and want to use them, and their preferences about privacy. Much of this work falls under the fields of psychology, user-interface design, and social sciences. This section is an introduction to these topics.
@@ VII-A Psychological privacy and ownership
People tend to have a deep sense of ownership of some things, both material and immaterial. In particular, people have a feeling of ownership toward information about themselves [186, 187]. Such feelings are, to some extent, detached from the material consequences of breaches in privacy. For example, even if publishing an audio recording revealing a secret intimate relationship would not have direct economic consequences, it can cause psychological damage and harm personal relationships. The effect of threats to privacy then necessarily becomes a question of psychology and social sciences. To qualitatively or quantitatively measure such effects, we also need subjective user tests.
Concurrently, people have well-established social rules regarding human-to-human privacy [25]. Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [188, 189], such that they will likely assume that devices apply human-like social rules to privacy. Observe that such social rules are related to human-like behavior and performance. That does not reveal whether they are applicable to the super-human performance that computers can possess, such as the ability to integrate information over massive databases and to retain accurate records of events long past. How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [190]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
@@ VII-B Perception and experience of privacy
Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [191]. Clearly, users do not like using services they perceive as threatening. Understanding how people perceive privacy is interesting in its own right, but such understanding is essential in the design of effective user interfaces.
Perception and experience of privacy can be approached in two alternative ways. We can make user studies where people interact with either machines or each other [192]. The distinction is important in the sense that human-to-human interaction relies on social rules that are well-established and developed over a long time. We can thus expect them to be stable over time, whereas human-computer interaction is continuously evolving as people learn more. Moreover, by using studies of human-to-human interaction, we can learn effects related to human-like performance but can probably not rely on observing effects related to super-human performance. Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [192].
Studies of human-computer interaction thus characterize the desired phenomenon directly. The compromise is, however, that people’s understanding of privacy with devices might not reflect the true level of privacy, and conclusions made based on the users’ opinions might then not reflect their true preferences and those preferences are moving targets. As people gain more experience with and as they learn more about technology, their attitudes, and perceptions of it change [192].
Despite these shortcomings, both types of experiments are essential for improving our understanding of privacy and for improving technology. For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [193, 17, 18, 194, 195] and multiple-bed patient rooms in hospitals have privacy-concerns [196]. Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [197], voice interfaces with unknown features cause fear in users, and reduces retention of services [198, 199, 200] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [201].
@@ VII-C User interface design
For usable privacy, it is important that information about privacy is readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [197, 202, 203, 199, 201, 204, 205, 206]. As a parallel to speech interfaces, visual interfaces can use lights or icons for monitoring and tactile interfaces for controlling the level of privacy. While sound can be used to monitor system status with sonification [207, 208], this is not in widespread use within signaling of privacy. It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [209]. To be effective, the sound should be perceivable when the user consciously wants to check the privacy level and changes in privacy level should evoke correspondingly and appropriately large changes in the soundscape that users consciously register such changes. Furthermore, user interfaces should give correct information about and allow controlling the privacy level [199, 15].
In any case, it is imperative that services are designed such that they reflect the true extent of privacy. Observe that it is likely possible to design systems that communicate an advanced level of privacy even when the system does not respect user privacy. In fact, service providers may have short-term incentives to follow such approaches to design as long as they improve overall user satisfaction. However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [210]. Through deception, such dark patterns lull users into believing they are safe when, in fact, they are abused. Therefore, good design of privacy should actively communicate and enable control of the true extent of privacy. Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [201].
@@ VII-D Privacy and trust
The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [192]. In particular, if users experience that a service is detrimental to their privacy or, more generally, to their well-being, they will obviously not trust it. We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [211].
While the meaning of trust may seem intuitively clear, it is however difficult to define exactly. A classic definition is [212]
@@
This definition highlights the psychological nature of trust [213]. From the current engineering perspective, this translates to the statistical confidence that the outcome of a process is useful. Here, usefulness refers to what the user, subjectively perceives as useful. Implicitly, this furthermore requires that we predict the outcome of that process in the form of a statistical distribution of possible outcomes and that the expectation over the possible outcomes is that the process is useful. Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [214, 215].
Observe the semantic difference between a system that is “benevolent” and one that is “useful”. The casual use of the concept of trust seems to require that the actor is indeed benevolent. However, it is possible to imagine a system with malicious objectives concerning Alice but produces side effects that consistently benefit Bob. Bob thus could potentially trust the system to act beneficially even if the system is not benevolent towards him. It is furthermore questionable whether we can assign an “intent” to a computational system due to its close association with psychology. The concept of benevolence is thus ambiguous in the context of speech technology.
It is useful and ethical to design speech services that are trustworthy. This means that systems should be such that it would be logical for users to trust them [216]. Speech technology should thus be designed to have useful outcomes with statistically significant confidence, especially with respect to privacy.
A practical approach for reducing the likelihood of negative outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [217]. In fig. 4, this can be implemented by applying layers of privacy protection in each successive connection and device as well as in their internal modules and components, as described in section V-E. Activation of an undesired functionality of the system or leaking private information would thus require that every subsequent module fails simultaneously, which is unlikely in most scenarios.

### v4

@@ VI Evaluating the performance of privacy-preserving methods
To evaluate the performance of any privacy-preserving methods, we need performance measures for both utility and privacy, corresponding respectively to the trusted and threat tasks (see fig. 2). With utility, we refer here to the usefulness of the service in its intended task. While the objective performance in both utility and privacy is a prerequisite for any good speech application, to the user, it is central that their experience with the service or application is good. Correspondingly, performance can always be measured on two dimensions, privacy and utility, as well as objective and subjective (see table IV). Performance with respect to objective and subjective privacy and utility is presented in the following subsections (sections VI-A, VI-B, VI-C and VI-D), but the subjective experience of privacy is a sufficiently large and independent topic that it deserves its own section (section VII). We further need to define how the objective measures are applied in systems design to ensure privacy (see section V-F).
The evaluation protocol of the VoicePrivacy challenge is a state-of-the-art, practical example of evaluation of both privacy and utility [tomashenko2022voiceprivacy]. It is based on direct measurement of identifiability using signal processing and machine learning methods. The evaluation objective in such empirical privacy is to determine the extent to which an attacker can re-identify an anonymized speaker or anonymized attribute(s) of a speaker (see fig. 3).
@@ VI-A Objective measures of privacy
A central issue with these metrics is that they are based on mathematical models of probability distributions, whereas speech is a natural signal that does not submit to exact formulas [nautsch20_interspeech]. To use these formulas, we must, therefore, first choose models to describe the distributions of observations and then estimate their parameters. Both steps lead to inaccuracies that dilute the evidentiary power of the proofs of privacy.
Measures commonly used to analyze privacy in speaker recognition, verification, and biometrics include [natgunanathan2016protection]:
@@
    The equal error rate (EER) is the proportion of misclassified samples when the threshold of acceptance is adjusted to give equal FJR and FAR [maouche20_interspeech, ISObiometric]. An increase in EER means that the attacker has made more errors, and privacy has improved.
@@
    The log-likelihood ratio (LLR), defined as the logarithm of the ratio of likelihoods $p(Y|H_{k})$ of the observation $Y$ conditioned with the hypotheses that either the sample is $H_{1}$, or is not $H_{0}$, the target speaker [nist2021SRE]
@@
    The application-independent log-likelihood-ratio cost function $C^{\min}_{llr}$ generalizes the EER by considering optimal thresholds over all possible prior probabilities and all possible error cost functions [maouche20_interspeech, brummer2006application].
@@
    The expected privacy disclosure from the Privacy ZEBRA framework [nautsch20_interspeech] quantifies the empirical cross-entropy (the leak magnitude) integrated over all possible prior likelihoods.
@@
    The worst-case disclosure also from the Privacy ZEBRA framework [nautsch20_interspeech] is the largest disclosure observed in a dataset.
@@
As mentioned above, empirical evaluation of privacy using EER has been applied, for example, in the VoicePrivacy challenge [tomashenko2022voiceprivacy]. Both EER and $C^{\min}_{llr}$ are widely used in any speaker recognition [KINNUNEN201012]. The worst-case disclosure in a quantized bottleneck in speaker recognition has been studied in [vali2024porcupine]. More comprehensive treatments of objective privacy metrics are provided in [nautsch2019preserving, noe2022towards].
@@ VI-B Subjective evaluation of privacy
In speaker recognition, computers are now on par or better than humans [hansen2015speaker] and are not fooled by voice mimicry (i.e., voice actors) even when humans struggle to do so [vestman2020methods, vestman2020voice]. We can thus expect that computers are better at evaluating privacy in well-defined tasks. However, there are at least two caveats. First, in identifying private information, human listeners can potentially use properties or features of speech signals other than those computers use. Subjective listeners can thus potentially discover new weaknesses that computers miss (though the opposite is also true). Second, in scenarios where humans are the potential attackers extracting information (as in threats 3, 5, and 7 in fig. 1), the true measure of privacy is quantified by subjective listening. Thus, it is often prudent to complement objective privacy metrics with subjective tests.
@@
As an example, in the VoicePrivacy 2020 challenge, subjective speaker verifiability was measured by asking subjects to rate, on a scale of 1 to 10110, the similarity between an anonymized target and a reference speaker, where the reference was either the same or a different speaker [tomashenko2022voiceprivacy]. If the distributions of similarity scores have significant differences, then the anonymization has failed. Alternatives to the 1 to 10110 Likert-scale include, for example, AB (are these samples the same person?) and ABX (which of AB is equal to X?) tests. For a comprehensive exposition of subjective evaluation methods, see [zacharov2018sensory].
@@ VI-C Objective performance in downstream tasks
Metrics used for quantifying utility depend on the application. In applications where a computer analyzes the speech signal, we quantify utility by measuring the accuracy of analysis. For example, in automatic speech recognition, where the computer extracts the text content of a speech signal, a commonly used measure is word error rate (WER), which is the proportion of correctly identified words [rohit2024asr]. By measuring the word error rate with and without privacy-preserving processing, we can determine how much privacy protections impede utility.
@@ VI-D Subjective performance in downstream tasks
If a human user is to listen to the output signal, then that user is the ultimate measure of performance [zacharov2018sensory]. For example, if we implement speech enhancement to reduce the perceived noise in a signal, then a user can subjectively evaluate how close the enhanced signal is to the noiseless original [vincent2018audio]. Such perceptual evaluation is best implemented using subjective listening test standards such as P.800 or MUSHRA [P800, MUSHRA]. Similarly, if privacy protections degrade a speech signal, we can readily quantify such degradations by listening tests.
Subjective listening tests are, however, expensive, time-consuming, and noisy. Individual subjects, especially those without training (known as naïve subjects), are typically unable to repeat a sequence of tests such that results would match. Subjects are also slow because they must listen to samples one at a time, at real-time speeds, and require time to answer between each sample. Additionally, a multitude of subjects need to complete the same test to gain statistical confidence in the results, and tests require oversight by trained personnel. Automated tests, known as objective measures of perceptual quality, are an alternative to such subjective tests. They estimate the outcome of subjective listening tests by computational means. For example, a classic objective measure known as PESQ estimates the outcome of a P.800 test [rix2001perceptual]. Later improvements include POLQA and ViSQOL [beerends2013perceptual, chinen2020visqol]. Such objective measures can be used to automate quality testing at a lower price and faster speed, so that repeated estimates with the same input give the same result. Observe that researchers should carefully review the specifications of such algorithms to determine whether they apply to the problem at hand to ensure reliable results.
@@ VII Perception and psychology of privacy
Much of the work of the speech community is focused on technical aspects of privacy-preserving algorithms. Yet speech devices are used by and designed for people, and they use people’s voice data. It is therefore essential to understand people’s perceptions of speech devices, including how they use and want to use them, as well as their preferences regarding privacy. Much of this work falls under the fields of psychology, user-interface design, and social sciences. This section is an introduction to these topics.
@@ VII-A Psychological privacy and ownership
People tend to have a deep sense of ownership of some things, both material and immaterial. In particular, people have a feeling of ownership toward information about themselves [cichy2014extending, dawkins2017psychological]. Such feelings are, to some extent, detached from the material consequences of breaches in privacy. For example, even if publishing an audio recording revealing a secret intimate relationship would not have direct economic consequences, it can cause psychological damage and harm personal relationships. The effect of threats to privacy then necessarily becomes a question of psychology and social sciences. To qualitatively or quantitatively measure such effects, we also need subjective user tests.
Concurrently, people have well-established social rules regarding human-to-human privacy [petronio2002boundaries]. Moreover, users also tend to anthropomorphize technology, that is, treat devices and services as if they were human [nass1993anthropomorphism, cornelius2021acceptance], such that they will likely assume that devices apply human-like social rules to privacy. Observe that such social rules are related to human-like behavior and performance. That does not reveal whether they apply to the super-human performance that computers can possess, such as the ability to integrate information over massive databases and to retain accurate records of events long past. How machines should then behave, especially with respect to super-human capabilities, is then not only a question of user-interface design but of moral psychology [thompson2021machine]: We need discussions on a societal level of how automated services should behave and how they are allowed to behave (see also section VIII).
@@ VII-B Perception and experience of privacy
Irrespective of whether a system actually preserves privacy or not, some users can perceive the system as threatening, and others can be oblivious to the threats it poses [pepino2020distrust]. Clearly, users do not like using services they perceive as threatening. Understanding how people perceive privacy is interesting in its own right, but such understanding is essential in the design of effective user interfaces.
Perception and experience of privacy can be approached in two alternative ways. We can make user studies where people interact with either machines or each other [leschanowsky2024review]. The distinction is important in the sense that human-to-human interaction relies on social rules that are well-established and developed over a long time. We can thus expect them to be stable over time, whereas human-computer interaction is continuously evolving as people learn more. Moreover, by using studies of human-to-human interaction, we can learn effects related to human-like performance, but can probably not rely on observing effects related to super-human performance. Studies of human-to-human interaction are, however, always a proxy if the actual target is to design human-to-computer interfaces [leschanowsky2024review].
Studies of human-computer interaction thus characterize the desired phenomenon directly. The compromise is, however, that people’s understanding of privacy with devices might not reflect the true level of privacy, and conclusions made based on the users’ opinions might then not reflect their true preferences and those preferences are moving targets. As people gain more experience with and as they learn more about technology, their attitudes and perceptions of it change [leschanowsky2024review].
Despite these shortcomings, both types of experiments are essential for improving our understanding of privacy and for improving technology. For example, human-to-human studies have revealed that people experience privacy differently in different acoustic environments; a noisy cafeteria can be better at masking sounds and defending against eavesdropping than a reverberant hallway [zarazaga19_interspeech, das2021thesis, zarazaga2022thesis, backstrom2021intuitive, Leschanowsky2020], and multiple-bed patient rooms in hospitals have privacy-concerns [reinten2017speech]. Similarly, human-machine studies have revealed that when a chatbot actively communicates choices related to privacy it improves users’ experience of privacy [bruggemeier2022perceptions], voice interfaces with unknown features cause fear in users, and reduces retention of services [yeasmin2020privacy, Lau2018, lutz2021privacy] and breaches in privacy are highly detrimental to the trust in services and reduce users’ willingness to use the services, but such trust can be rebuilt [Kurz2021].
@@ VII-C User interface design
For usable privacy, information about privacy must be readily provided, changes in the extent of privacy are promptly notified, and that users have control of the level of privacy [bruggemeier2022perceptions, so2021voice, lewis1998designing, Lau2018, Kurz2021, dhiya2023privacy, iachello2007end, sivaraman2018smart]. As a parallel to speech interfaces, visual interfaces can use lights or icons for monitoring and tactile interfaces for controlling the level of privacy. While sound can be used to monitor system status with sonification [hildebrandt2016continuous, iber2021auditory], this is not in widespread use within signaling of privacy. It is a compromise between filling the acoustic space with information and the tendency of the human auditory system to block out monotonous sounds [pulkki2015communication]. To be effective, the sound should be perceivable when the user consciously wants to check the privacy level, and changes in privacy level should evoke correspondingly and appropriately large changes in the soundscape so that users can consciously register the changes. Furthermore, user interfaces should provide accurate information about and enable control over the privacy level [Lau2018, backstrom2020privacy].
In any case, it is imperative that services are designed such that they reflect the true extent of privacy. Observe that it is likely possible to design systems that communicate an advanced level of privacy even when the system does not respect user privacy. In fact, service providers may have short-term incentives to follow such approaches to design as long as they improve overall user satisfaction. However, such approaches to design are known as dark patterns or deceptive design patterns, and they are considered unethical [willis2020deception]. Through deception, such dark patterns lull users into believing they are safe when, in fact, they are being abused. Therefore, good design of privacy should actively communicate and enable control of the true extent of privacy. Such design practice is not only ethical but also rewards service providers in the long term by improving retention of those services [Kurz2021].
@@ VII-D Privacy and trust
The level of privacy provided by a speech service is intimately linked to the level of trust users experience in that service [leschanowsky2024review]. In particular, if users experience that a service is detrimental to their privacy or, more generally, to their well-being, they will obviously not trust it. We thus discuss trust mainly from a psychological perspective, unlike in security research and cryptography, where trust has a specific and unambiguous mathematical definition [balsa2022cryptography].
While the meaning of trust may seem intuitively clear, it is, however, difficult to define exactly. A classic definition is [deutch1973resolution]
@@
This definition highlights the psychological nature of trust [simpson2007psychological]. From the current engineering perspective, this translates to the statistical confidence that the outcome of a process is useful. Here, usefulness refers to what the user subjectively perceives as useful. Implicitly, this furthermore requires that we predict the outcome of that process in the form of a statistical distribution of possible outcomes and that the expectation over the possible outcomes is that the process is useful. Another way of expressing this is to say that trust requires dimensions of competence, integrity, and benevolence [Chen2003, Xie2009].
Observe the semantic difference between a system that is “benevolent” and one that is “useful”. The casual use of the concept of trust seems to require that the actor is indeed benevolent. However, it is possible to imagine a system with malicious objectives concerning Alice, but that produces side effects that consistently benefit Bob. Bob thus could potentially trust the system to act beneficially even if it is not benevolent towards him. It is furthermore questionable whether we can assign an “intent” to a computational system due to its close association with psychology. The concept of benevolence is thus ambiguous in the context of speech technology.
It is useful and ethical to design speech services that are trustworthy. This means that systems should be such that it would be logical for users to trust them [hardin1996trustworthiness]. Speech technology should thus be designed to have useful outcomes with statistically significant confidence, especially with respect to privacy.
A practical approach for reducing the likelihood of adverse outcomes is to build multiple layers of checks and verifications, similar to multi-factor authentication [ometov2018multi]. In fig. 4, this can be implemented by applying layers of privacy protection in each successive connection and device, as well as in their internal modules and components, as described in section V-E. Activation of an undesired functionality of the system or leaking private information would thus require that every subsequent module fails simultaneously, which is unlikely in most scenarios.